40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 22 18:15:21 2010, kiwamu, Update, Locking, Green era: found a green beat note finally green_beatnote.jpg
    Reply  Fri Oct 22 21:28:27 2010, Aidan, Update, Locking, Green era: found a green beat note finally 
       Reply  Sun Oct 24 18:06:35 2010, kiwamu, Summary, Locking, setup for green beat green_beat_note.png
          Reply  Mon Oct 25 02:14:38 2010, Koji, Summary, Locking, setup for green beat 
Message ID: 3771     Entry time: Sun Oct 24 18:06:35 2010     In reply to: 3766     Reply to this: 3774
Author: kiwamu 
Type: Summary 
Category: Locking 
Subject: setup for green beat 

green_beat_note.png

(notes on signal level)

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

 

The expectation was calculated by using the following simple math.

                       SIGNAL = A x Z x G_RF x sqrt(P1 / 2) x sqrt (P2 / 2)

 where A is the responsibility of the PD and Z is the trans impedance gain. G_RF is a gain of the RF amplifier.

The laser powers of green beams, P1 and P2, are divided by 2 due to a beam splitter. 

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

 

ELOG V3.1.3-